# Campus5G: A Campus Scale Private 5G Open RAN Testbed

### v1

@@ Abstract.
copyright: none
@@ 1. Introduction
Mobile networks in the recent past have embraced disaggregation at different levels, driven by the need for cost-effective and flexible system architecture that accelerates innovation and enables new services. Complimentarily, they are headed towards data-driven and AI powered operation to unlock significant energy, operational and spectral efficiency gains. These trends are most prominently evident with Open Radio Access Network (RAN) (Wani et al., 2025), being standardized by the O-RAN Alliance (ora, [n. d.]a).
@@
It is in this context that private 5G mobile networks (Peterson et al., 2023) are seen as the early adopters of Open RAN. Unlike national scale public mobile networks, private networks are “local” scale deployments that target industrial (e.g., manufacturing, mining, warehouses) and enterprise (e.g., business parks, universities, hospitals) settings, transport hubs, venues and such (Kasujee and Rebbeck, 2023). Private mobile networks are appealing for these environments because they not only allow high degree of control and customization but also offer greater opportunity for innovation (e.g., edge AI applications (eri, 2024; Foukas and Radunovic, 2025)). Private 5G deployments are rapidly growing (pri, [n. d.]; bmw, 2022; bos, 2019; ara, [n. d.]; pri, 2024; uk-, 2025). Private 5G is a key use case for Open RAN (Martian et al., 2025). New breed of operators and vendors target private 5G with Open RAN as a default (tel, [n. d.]; neu, [n. d.]; jet, [n. d.]; Accelleran, [n. d.]).
@@ 2. Related Work
Arena (Bertizzolo et al., 2020) 24 Indoor Partially SDRs Open Source
Powder (pow, [n. d.]) 8 Outdoor Partially SDRs Open Source
Colosseum (Villa et al., 2024) 128 Emulated Partially SDRs Open Source
X5G (Villa et al., 2025) 8 Indoor Yes Commercial RUs
@@ 20 Indoor Yes Commercial RUs
Many testbeds currently exist to support O-RAN research. These testbeds by their nature are private (4G/5G) networks. Majority of them rely on USRP software defined radio (SDR) based radio frontends for the RUs, and are for the most part deployed indoors. Arena testbed (Bertizzolo et al., 2020) is an early example. NEC testbed (Salvat et al., 2023) is another. NSF PAWR platforms (paw, [n. d.]) also broadly fall under this category, though each with a different focus (e.g., mmWave in COSMOS (Chen et al., 2023), UAVs for AERPAW (Mushi et al., 2024), rural applications for ARA (Islam et al., 2025)). Of these, Powder (pow, [n. d.]) features an 8-node outdoor testbed with SDRs deployed across the campus of University of Utah. Similar SDR based Open RAN testbeds exist outside the US, including the Open Ireland testbed (ope, [n. d.]), testbeds in Romania and Norway (Martian et al., 2025) and various testbeds under Fed4Fire and follow on projects (fed, [n. d.]). Note that these SDR based testbeds do not fully support the O-RAN fronthaul interface between DU and RU based on the 3GPP 7.2x split that puts High-PHY on the DU side and Low-PHY on the RU side (Polese et al., 2023). As such, they are only partially compliant with the O-RAN standard (Martian et al., 2025) and consequently also are not compatible with O-RAN compliant commercial virtualized RAN (vRAN) solutions. Colosseum facility at Northeastern University (Villa et al., 2024) is also based on SDRs but relies instead on hardware based channel emulation, and so can support experimentation based on replaying real-world channel measurement traces.
In contrast to the above, other testbeds have emerged recently that use commercial RUs and are fully O-RAN compliant. The testbed at Microsoft Research, Cambridge (Bahl et al., 2023) is a building scale indoor deployment with 4 O-RAN compliant Foxconn RPQN-7800 RUs per floor across 5 floors, and uses an Intel FlexRAN based Layer 1 (L1), i.e., PHY (Intel, 2025a). X5G testbed at Northeastern is also an indoor testbed that consists of 8 Foxconn RPQN7801 RUs but instead uses NVDIA ARC based GPU accelerated L1. Recently, a nation wide Open RAN testbed deployment effort was reported from Brazil (Mauricio et al., 2025). While this testbed is large scale in its geographical span covering multiple interconnected sites across Brazil, each site has a small scale deployment with around handful of commercial RUs (from Foxconn, Benetel and VVDN), mostly indoor. Our Campus5G, while broadly in the same category as the above three testbeds, is deployed outdoors at a campus scale in a dense urban outdoor environment. It is currently based on Intel FlexRAN based L1 but is expected to also feature GPU accelerated L1 in the near future. Table 1 summarizes the comparison of Campus5G  with representative existing testbeds.
@@ 3.1. Radio Placement and Spectrum
To achieve our aim of blanket 5G campus wide coverage, we did not have unlimited flexibility for where we could deploy the RUs. Rather, we had to work within certain constraints placed by the heritage status of several of the university buildings and the feasibility of installing new fiber to buildings to support the O-RAN fronthaul requirements. This exercise resulted in identifying six university building sites where we could place the RUs.
@@
Given the above constraints, we leveraged our knowledge of the campus environment and footfall in different areas to come up with a radio network plan that would likely achieve blanket coverage across campus. This plan consisted of 20 RUs distributed across 6 building sites on campus along with the choice of power level, antenna type (omni or directional) and orientation for each of them. Fig. 1 (left) illustrates this plan overlaid on the campus map.
@@ 3.2. Pre-Deployment Coverage Map Estimation
Here we estimate the radio signal coverage map with our above outlined radio network plan. A coverage map (NVIDIA, 2025) associates signal quality metrics to each discrete bin on a 2D plane, enabling mobile operators to evaluate signal reception at arbitrary locations. We mainly focus on the received signal strength indicator (RSSI) metric. To achieve a proper trade-off between coverage map resolution and computational complexity, the estimated coverage maps are obtained at the $1$m×$\times$$1$m bin granularity within the campus area.
@@
We now estimate the RSSI coverage maps on the receiver plane at $1$m×$\times$$1$m bin granularity for three distinct cases: 1) switching on all the radios; 2) only switching on medium-power radios; and 3) only switching on low-power radios. Each of those coverage maps is obtained using the highest RSSI among all the transmitters computed at each bin. As shown in Fig. 2, when all the 20 radios are switched on, the estimated coverage map reflects blanket coverage across the campus (between $-65$ and $-50$ dBm), with particularly strong coverage in the center of the campus area. When only medium-power radios are activated, most regions across the campus still experience satisfactory signal coverage (above $-80$ dBm) as illustrated in Figure 2. But certain areas highlighted by red circles show significant coverage deterioration compared with the scenario with all radios active. This result validates our radio network planning strategy, emphasizing the essential role of low-power directional radios in significantly enhancing signal quality in certain parts of the campus.
@@ 4.1. Architecture and Design Choices
From a testbed design perspective, besides the number of radios and their placement covered in the previous section, we have few further design decisions to make: (1) choosing an O-RAN deployment scenario; (2) fronthaul design; (3) handling time synchronization. For (1), we opt for the O-RAN deployment scenario “A” (ora, 2023) where CU, DU and near real-time (RT) RIC are collocated at an edge cloud while RUs are realized as physical network functions that communicate with DU over the O-RAN fronthaul. This scenario is most suitable for a campus scale deployment as all RAN related processing and software can be consolidated in one edge cloud data center location, thereby giving ample flexibility to decide on the different L1 processing options (GPU accelerated or not), vRAN alternatives (commercial or open-source), RICs (standard or research) and corresponding set of Apps.
@@ 4.2. Testbed Components
Platform. The role of the platform is to provide a software layer atop the hardware, for all the upper layers to run on. The platform is a key enabler of flexibility, providing a base for running a diverse set of possible RANs, RICs and Apps. For the base layer, i.e., the operating system (OS), we chose Ubuntu (Canonical, [n. d.]), optimizing it to support real-time processing – a necessity for the RAN functions. Atop the OS, Ansible (Red Hat, [n. d.]) was used to provision basic utilities such as networking, some drivers, and the Kubernetes orchestrator (Foundation, [n. d.]b). Kubernetes is the main component of the platform. All of the workloads intended to be deployed (RANs, RICs and Apps) were containerized and deployed atop it. To manage this orchestration, we made extensive use of Helm charts (Foundation, [n. d.]a), which allowed ‘one-click’ deployment of a specified RAN setup.
@@
We use srsRAN as the default vRAN software in our deployment and use a split CU-DU architecture with a single CU mapped to 20 DUs. This setup eases mobility management, specifically enabling intra-CU inter-DU handovers and ensuring seamless user transitions between cells. For the core network, we currently use an Open5GS (Open5GS, 2025) based cloud-hosted solution. Our testbed supports access from commodity devices through custom eSIM and physical SIM cards.
RICs and Apps. Our platform supports the deployment of standard O-RAN E2 (O-RAN, 2022) based RICs, enabling real-time network optimization through intelligent control loops via xApps. Specifically, we have successfully integrated O-RAN SC RIC (ora, [n. d.]b) and FlexRIC (Schmidt et al., 2021) along with standard applications such as the KPIMON xApp, facilitating basic monitoring and control functionalities. Given the rising interest in non-standard RICs, which offer more granular control over RAN operations, we have additionally integrated two research RIC solutions: EdgeRIC (Ko et al., 2024) and Microsoft Janus/jbpf (Foukas et al., 2023; jbp, [n. d.]). Our testbed also features advanced AI/ML based Open RAN Apps (e.g., SpotLight (Sun et al., 2024)).
@@ 4.3. RAN Configuration
However, we found that not all vendor defaults were viable in our environment, and several needed to be manually tuned or overridden. For instance, we deployed 20 radios overall from two different RU vendors, VVDN and Benetel, each supporting different feature subsets. VVDN RUs, for example, support only a fixed TDD configuration and short PRACH format, while Benetel supports seven TDD configurations along with both short and long PRACH formats. Additionally, each RU vendor specifies its own fronthaul delay profile (O-RAN, 2024), but these defaults often do not reflect the real-world deployments. In one case, migrating a VVDN RU from a short DAC cable based lab testbed to a rooftop site over fiber introduced additional propagation delay, pushing eCPRI timestamps outside the DU’s expected window. This resulted in packet loss until we manually adjusted the delay profile to align with the physical link characteristics.
With 20 radios deployed across our campus, most locations observe coverage from 2–3 cells simultaneously. Such a deployment density required careful PRACH planning (Zhang, 2017). We allocated a unique $prach\_root\_sequence\_index$ to each cell site to prevent random access (RACH) collisions or failures. These configurations were sufficient to bring up a stable and functional RAN across our deployment. However, they are not yet fully optimized. Many additional parameters require further tuning and per-site calibration to reach optimal performance.
With our current deployment and default TDD configuration, we have achieved max downlink (uplink) throughput of 650 Mbps (around 80Mbps). This is lower than the absolute maximum possible throughput of 1.1Gbps with 100MHz and 4x4 MIMO, suggesting room for optimization through configuration parameter tuning. To put these numbers in context, what we already achieve is more than 3.5 times the max downlink throughput measured with public 5G mobile networks deployed in our campus area.
@@ 5.2. Measurement Based Coverage Maps
We conducted a  15 Km walking test across the campus to collect measurements using a commodity 5G smartphone, specifically Nothing Phone 2 (np2, [n. d.]), including GPS locations and UE reported metrics (RSRP, RSRQ and SINR). Around 11K samples in the form of [GPS location, reported metrics] are collected. To align the metrics obtained from Sionna and field measurements, we compute RSSI from the measured data based on the following relationship (3rd Generation Partnership Project (3GPP), 2023):
@@
Figure 3. Measurement based RSSI (left) and SINR (right) coverage maps.
@@ 8. Conclusions
In this paper, we have captured our journey planning, deploying and measuring Campus5G, the first-of-its kind campus scale O-RAN testbed. Urban, outdoor and university campus setting that our deployment represents is a perfect blend for innovation on future wireless access in a real-world environment. We reflect on our experience deploying Campus5G and share the key lessons. We also outline a number of research opportunities that came to the surface through our deployment experience.
@@ Appendix A Overview of Interpolation Methods
  • Machine-learning-based. In this category, location coordinates (and any covariates) form a feature vector $\mathbf{x}$, and then a flexible function $\hat{Z}(\mathbf{s}_{0})=f(\mathbf{x}_{0};\,\theta)$ is learned by minimising an interpolation loss (Mohammadjafari et al., 2020), where $\theta$ denotes unknown parameters. We consider two effective ML-based algorithms that are widely used, namely Random Forest (RF) and XGBoost. Both algorithms scale well to large spatial datasets and require much lighter computation than the neural network based methods.

### v2

@@ Abstract.
Open RAN, Private 5G, Testbeds
copyright: noneccs: Networks Mobile networksccs: Networks Network experimentation
@@ 1. Introduction
Mobile networks in the recent past have embraced disaggregation at different levels, driven by the need for cost-effective and flexible system architecture that accelerates innovation and enables new services. Complimentarily, they are headed towards data-driven and AI powered operation to unlock significant energy, operational and spectral efficiency gains. These trends are most prominently evident with Open Radio Access Network (RAN) (Wani et al., 2025), being standardized by the O-RAN Alliance (O-RAN ALLIANCE e.V, [n. d.]).
@@
It is in this context that private 5G mobile networks (Peterson et al., 2023) are seen as the early adopters of Open RAN. Unlike national scale public mobile networks, private networks are “local” scale deployments that target industrial (e.g., manufacturing, mining, warehouses) and enterprise (e.g., business parks, universities, hospitals) settings, transport hubs, venues and such (Kasujee and Rebbeck, 2023). Private mobile networks are appealing for these environments because they not only allow high degree of control and customization but also offer greater opportunity for innovation (e.g., edge AI applications (eri, 2024; Foukas and Radunovic, 2025; McLean et al., 2022)). Private 5G deployments are rapidly growing (Observatory, [n. d.]; bmw, 2022; bos, 2019; Labs, [n. d.]; pri, 2024; uk-, 2025). Private 5G is a key use case for Open RAN (Martian et al., 2025). New breed of operators and vendors target private 5G with Open RAN as a default (Limited, [n. d.]b; neu, [n. d.]; jet, [n. d.]; Accelleran, [n. d.]).
@@ 2. Related Work
[…]
@@
Indoor and
Outdoor
Yes Commercial RUs Commercial
[…]
Indoor and
Outdoor
Yes Commercial RUs Commercial
[…]
@@
Many testbeds currently exist to support O-RAN research. These testbeds by their nature are private (4G/5G) networks. Majority of them rely on USRP software defined radio (SDR) based radio frontends for the RUs, and are for the most part deployed indoors. Arena testbed (Bertizzolo et al., 2020) is an early example. NEC testbed (Salvat et al., 2023) is another. NSF PAWR platforms (for Advanced Wireless Research Project Office, [n. d.]) also broadly fall under this category, though each with a different focus (e.g., mmWave in COSMOS (Chen et al., 2023), UAVs for AERPAW (Mushi et al., 2024), rural applications for ARA (Islam et al., 2025)). Of these, Powder (Breen et al., 2021) features an 8-node outdoor testbed with SDRs deployed across the campus of University of Utah. Similar SDR based Open RAN testbeds exist outside the US, including the Open Ireland testbed (ope, [n. d.]), testbeds in Romania and Norway (Martian et al., 2025) and various testbeds under Fed4Fire and follow on projects (Consortium, [n. d.]). Note that these SDR based testbeds do not fully support the O-RAN fronthaul interface between DU and RU based on the 3GPP 7.2x split that puts High-PHY on the DU side and Low-PHY on the RU side (Polese et al., 2023). As such, they are only partially compliant with the O-RAN standard (Martian et al., 2025) and consequently also are not compatible with O-RAN compliant commercial virtualized RAN (vRAN) solutions. Colosseum facility at Northeastern University (Villa et al., 2024) is also based on SDRs but relies instead on hardware based channel emulation, and so can support experimentation based on replaying real-world channel measurement traces.
In contrast to the above, other testbeds have emerged recently that use commercial RUs and are fully O-RAN compliant. The testbed at Microsoft Research, Cambridge (Bahl et al., 2023) is a building scale indoor deployment with 4 O-RAN compliant Foxconn RPQN-7800 RUs per floor across 5 floors, and uses an Intel FlexRAN based Layer 1 (L1), i.e., PHY (Intel, 2025a). X5G testbed at Northeastern is also an indoor testbed that consists of 8 Foxconn RPQN7801 RUs but instead uses NVDIA ARC based GPU accelerated L1. Recently, a nation wide Open RAN testbed deployment effort was reported from Brazil (Mauricio et al., 2025). While this testbed is large scale in its geographical span covering multiple interconnected sites across Brazil, each site has a small scale deployment with around handful of commercial RUs (from Foxconn, Benetel and VVDN), mostly indoor. The FCT O-RAN testbed at Singapore Institute of Technology (Rezende et al., 2025) mixes indoor (3 Foxconn RPQN780) and outdoor (2 Benetel RAN650) RUs with a FlexRAN L1 and Radisys CU / DU. Our Campus5G, while broadly in the same category as the above three testbeds, is deployed outdoors at a campus scale in a dense urban outdoor environment. It is currently based on Intel FlexRAN based L1 but is expected to also feature GPU accelerated L1 in the near future. Table 1 summarizes the comparison of Campus5G  with representative existing testbeds.
@@ 3.1. Radio Placement and Spectrum
The goal of the planning process was to find a suitable placement of the radios that would provide both blanket coverage across the entire campus and surrounding areas, as well as sufficient capacity for known high-footfall areas where we expected larger numbers of users to congregate or traverse on a regular basis. We also preferred to over-provision rather than the opposite, as over-provisioning allows us to slice the network on a radio basis; for example one slice can serve our production users while others can be used for experimental and integration purposes.
An additional constraint was that we did not have unlimited flexibility for where we could deploy the RUs. Rather, we had to work within certain constraints placed by the heritage status of several of the university buildings and the feasibility of installing new fiber to buildings to support the O-RAN fronthaul requirements. This exercise resulted in identifying six university building sites where we could place the RUs.
@@
Given the above constraints, we leveraged our knowledge of the campus environment and footfall in different areas to come up with a radio network plan that would achieve blanket coverage across campus, while ensuring sufficient capacity in all areas. Moreover, we optimized for reducing licensing cost where possible, by using medium-power licenses to reach most of the campus, with the remaining areas served by low-power licenses. This plan consisted of 20 RUs distributed across 6 building sites on campus along with the choice of power level, antenna type (omni or directional) and orientation for each of them. Fig. 1 (left) illustrates this plan overlaid on the campus map.
@@ 3.2. Pre-Deployment Coverage Map Estimation
Here we estimate the radio signal coverage map with our above outlined radio network plan. A coverage map (NVIDIA, 2025) associates signal quality metrics to each discrete bin on a 2D plane, enabling mobile operators to evaluate signal reception at arbitrary locations. We mainly focus on the received signal strength indicator (RSSI) metric. To achieve a proper trade-off between coverage map resolution and computational complexity, the estimated coverage maps are obtained at the $\times11m$ bin granularity within the campus area.
@@
We now estimate the RSSI coverage maps on the receiver plane at $\times11m$ bin granularity for three distinct cases: 1) switching on all the radios; 2) only switching on medium-power radios; and 3) only switching on low-power radios. Each of those coverage maps is obtained using the highest RSSI among all the transmitters computed at each bin. As shown in Fig. 2, when all the 20 radios are switched on, the estimated coverage map reflects blanket coverage across the campus (between 65-65 and 50-50 dBm), with particularly strong coverage in the center of the campus area. When only medium-power radios are activated, most regions across the campus still experience satisfactory signal coverage (above 80-80 dBm) as illustrated in Figure 2. But certain areas highlighted by red circles show significant coverage deterioration compared with the scenario with all radios active. This result validates our radio network planning strategy, emphasizing the essential role of low-power directional radios in significantly enhancing signal quality in certain parts of the campus.
@@ 4.1. Architecture and Design Choices
From a testbed design perspective, besides the number of radios and their placement covered in the previous section, we have few further design decisions to make: (1) choosing an O-RAN deployment scenario; (2) fronthaul design; (3) handling time synchronization. For (1), we opt for the O-RAN deployment scenario “A” (ora, 2023) where CU, DU and near real-time (RT) RIC are collocated at an edge cloud while RUs are realized as physical network functions that communicate with DU over the O-RAN fronthaul.
@@ 4.2. Testbed Components
Platform. The role of the platform is to provide a software layer atop the hardware, for all the upper layers to run on. The platform is a key enabler of flexibility, providing a base for running a diverse set of possible RANs, RICs and Apps. For the base layer, i.e., the operating system (OS), we chose Ubuntu (Canonical, [n. d.]), optimizing it to support real-time processing – a necessity for the RAN functions. Atop the OS, Ansible (Inc., [n. d.]) was used to provision basic utilities such as networking, some drivers, and the Kubernetes orchestrator (Foundation, [n. d.]b). Kubernetes is the main component of the platform. All of the workloads intended to be deployed (RANs, RICs and Apps) were containerized and deployed atop it. To manage this orchestration, we made extensive use of Helm charts (Foundation, [n. d.]a), which allowed ‘one-click’ deployment of a specified RAN setup.
@@
We use srsRAN as the default vRAN software in our deployment and use a split CU-DU architecture with a single CU mapped to 20 DUs (one per RU). This setup eases mobility management, specifically enabling intra-CU inter-DU handovers and ensuring seamless user transitions between cells. For the core network, we currently use an Open5GS (Open5GS, 2025) based cloud-hosted solution. Our testbed supports access from commodity devices through custom eSIM and physical SIM cards.
RICs and Apps. Our platform supports the deployment of standard O-RAN E2 (O-RAN, 2022) based RICs, enabling real-time network optimization through intelligent control loops via xApps. Specifically, we have successfully integrated O-RAN SC RIC (Software Community, [n. d.]) and FlexRIC (Schmidt et al., 2021) along with standard applications such as the KPIMON xApp, facilitating basic monitoring and control functionalities. Given the rising interest in non-standard RICs, which offer more granular control over RAN operations, we have additionally integrated two research RIC solutions: EdgeRIC (Ko et al., 2024) and Microsoft Janus/jbpf (Foukas et al., 2023; Microsoft, [n. d.]). Our testbed also features advanced AI/ML based Open RAN Apps (e.g., SpotLight (Sun et al., 2024)).
@@ 4.3. RAN Configuration
However, we found that not all vendor defaults were viable in our environment, and several needed to be manually tuned or overridden. For instance, we deployed 20 radios overall from two different RU vendors, VVDN and Benetel, each supporting different feature subsets. VVDN RUs, for example, support only a fixed TDD configuration and short PRACH format, while Benetel supports seven TDD configurations along with both short and long PRACH formats. Additionally, each RU vendor specifies its own fronthaul delay profile (O-RAN, 2024), but these defaults often do not reflect the real-world deployments. In one case, migrating a VVDN RU from a lab testbed using short direct-attach-copper cables to a rooftop site over fiber introduced additional propagation delay, pushing eCPRI timestamps outside the DU’s expected window. This resulted in packet loss until we manually adjusted the delay profile to align with the physical link characteristics.
With 20 radios deployed across our campus, most locations observe coverage from 2–3 cells simultaneously. Such a deployment density required careful PRACH planning (Zhang, 2017). We allocated a unique prach_root_sequence_indexprach\_root\_sequence\_index to each cell site to prevent random access (RACH) collisions or failures. These configurations were sufficient to bring up a stable and functional RAN across our deployment. However, they are not yet fully optimized. Many additional parameters require further tuning and per-site calibration to reach optimal performance.
With our current deployment and default TDD configuration, we have achieved max downlink (uplink) throughput of 800 Mbps (around 80Mbps). This is lower than the absolute maximum possible throughput of 1.4Gbps with 100MHz and 4x4 MIMO, suggesting room for optimization through configuration parameter tuning. To put these numbers in context, what we already achieve is more than 3.5 times the max downlink throughput measured with public 5G mobile networks deployed in our campus area.
@@ 5.2. Measurement Based Coverage Maps
We conducted a  15 Km walking test across the campus to collect measurements using a commodity 5G smartphone, specifically Nothing Phone 2 (Limited, [n. d.]a), including GPS locations and UE reported metrics (RSRP, RSRQ and SINR). Around 11K samples in the form of [GPS location, reported metrics] are collected. To align the metrics obtained from Sionna and field measurements, we compute RSSI from the measured data based on the following relationship (3rd Generation Partnership Project (3GPP), 2023):
@@
Figure 3. Measurement based RSSI coverage (left) and SINR (right) maps.
@@ 8. Conclusions
In this paper, we have captured our journey planning, deploying and measuring Campus5G, the first-of-its kind campus scale O-RAN testbed. Urban, outdoor and university campus setting that our deployment represents is a perfect blend for innovation on future wireless access in a real-world environment. We have reflected on our experience deploying Campus5G and shared the key lessons. We have also outlined a number of research opportunities that came to the surface through our deployment experience.
Acknowledgements.
This work was mainly supported by a project funded by the UK Department for Science, Innovation and Technology (DSIT). The work of T. Wang and M. Marina was also supported by the UKRI/EPSRC grant UKRI860.
@@ Appendix A Overview of Interpolation Methods
  • Machine-learning-based. In this category, location coordinates (and any covariates) form a feature vector $\mathbf{x}$, and then a flexible function $\hat{Z}(\mathbf{s}_{0})=f(\mathbf{x}_{0};\,\theta)$ is learned by minimizing an interpolation loss (Mohammadjafari et al., 2020), where $\theta$ denotes unknown parameters. We consider two effective ML-based algorithms that are widely used, namely Random Forest (RF) and XGBoost. Both algorithms scale well to large spatial datasets and require much lighter computation than the neural network based methods.